OpenStudy (anonymous):

true or false For every line segment there is exactly one midpoint. the answer is true but i have to draw a sketch and i need help with that

OpenStudy (anonymous):

its true and you can find the coordination of the midpint x(mid) = (x1 + x2)/2 y(mid) = (y1 + y2)/2

OpenStudy (anonymous):

wait what am i supposed to do with that

OpenStudy (anonymous):

you can find the midpoint

OpenStudy (anonymous):

i can find the midpoint with that?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

look make any line segment with end point and begining point and apply these equation to find the midpoint

OpenStudy (anonymous):



OpenStudy (anonymous):

is it clear ??

OpenStudy (anonymous):

i dont understanmd the whole xy thing